A neon-dioxygen mixture contains 70.6 g dioxygen and 167.5 g neon. If pressure of the mixture of gases in the cylinder is 25 bar. What is the partial pressure of dioxygen and neon in the mixture ? Number of moles of dioxygen

Grade:12

6 Answers

Gaurav
9 years ago
Hello Student
no2 = 70.6/32 = 2.21
nNe = 167.5/20 = 8.375
Xo2 = 2.21/(2,21+8.375) =0 .21
XNe = 0.79
Partial Pressure of a Gas = Mole Fraction
Partial pressure of Oxygen = 0.21
Oartial pressure of Neon = 0.79
Rajinder Bhat
6 years ago
Partial partial pressure is equal to mole fraction x total pressure of the mixtureFind the mole fraction as given w e i g h t upon molar mass
Mohd Aabshar
6 years ago
Partial pressure = mole fraction of given gas x pure pressure.Mole fraction = Number of moles of given gas / total number of moles.Mole of dioxygen in mixture = given mass of dioxygen / molecular mass of dioxygen = 70.6 g/ 32 g = 2.2 molesMoles of Neon in mixture = given mass of neon / atomic mass of neon = 167.5 g / 20 = 8.3 moles.total number of moles = 2.2 + 8.3 = 10.5 molemole fraction of O2 = 2.2 / 10.5 = 0.2Mole fraction of neon = 8.3/ 10.5 = 0.8partial pressure of di oxygen = 0.2 × 25 = 5 barPartial pressure of neon = × 25 = 20 bar
Ronny
6 years ago
According to Dalton's Law Of Partial Pressure
 
P(x)= P X1   ; [X1 is Mole fraction ]
 
X1 of Ne = n(Ne)/( n(Ne)+n(O2) )
                 =       167.5/20.        
                    167.5/20+70.6/32
                 = 0.791494
 
P(x) = P + X1
         = (0.083 × 25) + 0.791494
         = 2.866494
                  
VakulKeshav
5 years ago
Number of moles of O2 = 70.6/32 = 2.21 mol
Number of moles of neon = 167.5/20 = 8.375 mol
Mol fraction of dioxygen = 2.21/2.21+8.375 = 0.21
Mol fraction of neon = 1- 0.21 = 0.79
Partial pressure of gas = Mol fraction × total pressure
Partial pressure of oxygen = 0.21 × 25 = 5.24 bar
Partial pressure of neon gas = 0.79 × 25 = 19.75 bar
 
 
ankit singh
3 years ago

 

The mass of dioxygen is 70.6 g.
 
The molar mass of dioxygen is 32 g/mol.
 
The number of moles of dioxygen
 
= 32 g/mol  70.6 g = 2.21 mol 
 
The mass of neon is 167.5 g.
 
The molar mass of neon is 20.2 g/mol.
 
The number of moles of neon
 
= 20.2 g/mol  167.5 g = 8.29 mol 
 
The total number of moles of neon and dioxygen = 2.21 + 8.29= 10.49 
 
The mole fraction of neon = 10.49 8.29 = 0.789 
 
Total pressure in the cylinder is 25 bar.
 
The partial pressure of neon = 0.789 × 25 bar= 19.7 bar
 
The partial pressure of dioxygen = 25 bar  19.7 bar = 5.3 bar